The area of a trapezoid is 63 square inches. One base is 9 in. long and the other is 10 in long. Find the height of the trapezoid.
Answer by checkley71(8403)

63=H(9+10)/2 OR 63=H*19/2 OR 63=9.5H OR H=63/9.5 OR H=6.6315789
PROOF 63=6.6315789(9.5 OR 63=63